1 Singular systems, state feedback problem

By Marija Dodig.

In this talk, the strict equivalence invariants by state feedback for singular systems are studied. As the main result we give the necessary and sufficient conditions under which there exists a state feedback such that the resulting system has prescribed pole structure as well as row and column minimal indices. This result presents a generalization of previous results of state feedback action on singular systems.